# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v1

@@ 3.2 Policy Model Mis-specification
In this case, RLHF has a structural advantage: it can recover the exact reward and then compute the best possible policy within $\Pi$. In contrast, DPO bypasses reward modeling and directly learns a policy from preferences, which may lead to sub-optimal behavior due to mismatches between preference-based objectives and reward-based value functions. The following proposition provides a concrete example where DPO fails to recover the best achievable policy, even under exact optimization. Proof deferred to Section C.3.
@@
Furthermore, we show that online DPO cannot close this gap, even when equipped with PILAF sampler. A numerical proof is deferred to Section C.8.
@@ 3.3 Reward Model Mis-specification
We now consider the setting where the ground-truth reward function $r^{\star}$ is not realizable by the reward model class $\mathcal{F}$, while the optimal policy $\pi^{\star}$ lies within the policy class $\Pi$. This setting is studied for completeness. As discussed in Swamy et al. (2024), two-stage RLHF can only lose information during reward learning, which will be highlighted under reward model mis-specification.
@@
In this setting, RLHF is vulnerable to reward mis-specification: the learned mis-specified reward model $r_{\textup{RLHF}}$ could significantly deviate from the ground-truth reward $r^{\star}$, causing the subsequent policy optimization to yield a sub-optimal solution even though $\pi^{\star}\in\Pi$. Conversely, DPO has a clear advantage: it can directly fit a policy to the observed preference data and thus recover $\pi^{\star}$ without incurring reward modeling error. Proof deferred to Section C.4.
@@ 3.4 Double Model Mis-specification
To enable a fine-grained comparison between RLHF and DPO under this double mis-specified regime, we introduce the surrogate reward model class induced by the policy class as $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\theta\in\mathbb{R}^{d_{P}},\hat{r}_{\theta}(y)=\beta\log\frac{\pi_{\theta}(y)}{\pi_{{\mathsf{ref}}}(y)},\forall y\in\mathcal{Y}\}$. Pairwise preferences depend only on reward differences, so reward functions $r,r^{\prime}:\mathcal{Y}\to\mathbb{R}$ are equivalent if they differ by a constant. We compare the expressiveness of the original reward model class $\mathcal{F}$ and the surrogate class $\mathcal{F}_{\Pi}$, modulo constant shifts, and analyze three representative regimes characterizing their relative capacities:
@@
For Condition 5, we focus on the scenario when the reward model class and policy model class are isomorphic—meaning there exists a shared parameterization or a deterministic mapping between rewards and policies. This structure allows us to directly compare RLHF and DPO when both operate under the same representational constraints, and to investigate whether bypassing reward modeling, as in DPO, provides any advantage. Then motivated by Nika et al. (2024), we can split the performance of RLHF and DPO as (see detailed calculations in Section C.12):
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{RLHF}}}_{r^{\star}}$ $=\underset{y\sim\pi_{\textup{RLHF}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{RLHF}}(y)-\beta\log\pi^{\star}(y)\right]\;,$
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\textup{DPO}}}_{r^{\star}}$ $=\underset{y\sim\pi_{\textup{DPO}}}{\mathbb{E}}\left[\beta\log\pi_{\textup{DPO}}(y)-\beta\log\pi^{\star}(y)\right]\;.$
In two-stage RLHF, reward learning is decoupled from the current policy, and thus lacks access to its distributional information; while DPO can mitigate this limitation through online sampling. Therefore, two-stage RLHF under Condition 5 is comparable to offline DPO, but could underperform online DPO. Proofs deferred to Sections C.5 and C.9.
@@
On the other hand, under Conditions 6 and 7, either method may outperform the other depending on the environment. Proofs deferred to Sections C.6 and C.7.
@@
This result implies a general principle: to measure the performance gap, we only need to measure the quality gap between the (surrogate) reward models.
As revealed in Appendix C of Ouyang et al. (2022) and Section 3.3 of Swamy et al. (2025), it is uncommon to deploy a reward model with greater scale than the policy model. To ensure practical relevance, we focus on the regime where $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, and pose the following relevant question:
@@
As an answer to this question, we can show that in the context of preference-based policy learning, the reward model quality can be measured using an $\ell_{2}$ distance of pairwise difference: $\mathbb{E}_{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]^{2}$, where $\pi_{\theta^{\star}(r_{\phi})}:=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\phi}}^{\pi}$. The detailed calculations and further discussions are deferred to Appendix B. In Section 4, we adopt a data-dependent proxy for this notion: the empirical $\ell_{2}$ distance over observed preference pairs (see Definition 2 in Section C.10, also in Zhu et al. (2023)). Using this metric, we establish a separation result that demonstrates the statistical advantage of explicitly learning reward functions.
4 Approximate Optimization: Implicit Representation Gap Induced by Statistical Efficiency in Sparse Recovery
Under more realisitc finite-sample regime, statistical efficiency can also induce performance gap. With limited preference data, we are not able to directly compute $r_{\textup{RLHF}}$, $\pi_{\textup{RLHF}}$ and $\pi_{\textup{DPO}}$, and thus obtain weaker solutions due to estimation error. This scenario can be viewed as inducing an implicit representation gap, whose effects have been widely discussed in Section 3. Motivated by Nika et al. (2024) and still assuming $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, we can split the performance of two-stage RLHF as:
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}\leq\underbrace{\left(V^{\pi_{\theta^{\star}(r_{\phi})}}_{r_{\phi}}-V^{\pi_{\theta}}_{r_{\phi}}\right)}_{\textup{policy estimation error}}+\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-r_{\phi}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[r_{\phi}(y)-r^{\star}(y)\right]}_{\textup{implicit reward representation gap}}\;,$
and for DPO we have (see detailed calculations in Section C.13):
$V^{\pi^{\star}}_{r^{\star}}-V^{\pi_{\theta}}_{r^{\star}}$ $\leq\underbrace{\underset{y\sim\pi^{\star}}{\mathbb{E}}\left[r^{\star}(y)-\hat{r}_{\theta}(y)\right]+\underset{y\sim\pi_{\theta}}{\mathbb{E}}\left[\hat{r}_{\theta}(y)-r^{\star}(y)\right]}_{\textup{ implicit reward representation gap}}\;.$
We now compare the statistical efficiencies of RLHF and DPO, focusing on how their sample complexities differ and influence the implicit reward representation gap. We will further improve the analysis of Nika et al. (2024); Swamy et al. (2025), to prove that when the reward model and the policy model share the same model size, i.e. $d_{P}=d_{R}$ and even under realizability condition, two-stage RLHF is still more sample-efficient than DPO.
@@
Difference in token-level parameterization. The common reward model shares the same architecture with LM but replaces the last layer with a linear head, i.e., it takes the whole prompt-response pair as the input and predicts one value. Therefore, if we view the last-layer hidden state as the feature vector, it is natural to assume the reward model to be parameterized as a linear MDP model [^1] It is also common to assume the reward model to be a linear bandit model (Zhu et al., 2023), while the stronger linear MDP model assumption here is for fair comparison with the following policy model.: $r_{\theta}(y)=\beta\sum_{t=0}^{N}\theta_{t}^{\top}\psi(y_{0\ldots t})$, where $\theta,\psi(y)\in\mathbb{R}^{d}$, and $N=|y|-1$. While for the policy model, one needs to go through the softmax results of all tokens and multiply them [^2] Our parameterization assumption on the token-level policy model is different from Razin et al. (2025a), which utilizes a form of token matrix, since we intend to ensure that $d_{P}=d_{R}$.:
$\pi_{\theta}(y)=\prod_{t=0}^{N}\pi_{\theta}(y_{t}|y_{0\ldots t-1})=\prod_{t=0}^{N}\frac{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp(\theta_{t}^{\top}\psi(y_{0\ldots t}))}{\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(\theta_{t}^{\top}\psi(y_{0\ldots t-1},s))}\;.$
@@
Let the ground truth reward be $r^{\star}(y)=\beta\sum_{t=0}^{N}(\theta_{t}^{\star})^{\top}\psi(y_{0\ldots t})$. Then the optimal solution for the reward model is $\theta_{t}=\theta^{\star}_{t}$, $\forall t\in\mathbb{Z}_{\geq 0}$. And the optimal solution for the policy model is given by Rafailov et al. (2024). Here we interpret the result in a slightly different way to separate $\pi_{{\mathsf{ref}}}$ and the $q$ function (see detailed calculations in Section C.13):
@@
$\pi^{\star}(y_{t}|y_{0\ldots t-1})=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q(y_{t}|y_{0\ldots t-1})-q(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$ (2)
where the $q$ function is determined in a recursive way:
$q(y_{t}|y_{0\ldots t-1})=\begin{cases}\beta\log\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t})\exp(q(s|y_{0\ldots t})/\beta)&y_{t}\text{ is not the terminal token;}\\ r^{\star}(y_{0\ldots t})&y_{t}\text{ is the terminal token.}\end{cases}$
@@
Benefiting from the token-level $q$ function, models trained in this way can simulate a process reward model to provide fine-grained information (Yuan et al., 2024; Cui et al., 2025; Shi et al., 2024; Xu et al., 2025). However, simultaneously, learning the $q$ function sacrifices statistical efficiency due to the need to model the complicated structure. Next, we will present a concrete example to illustrate the statistical gap between pure reward learning and surrogate reward learning.
@@
Dual-token Sparse Prediction (DTSP).  Under log-linear policy class, the policy model is required to sequentially output two tokens, $y$ and $\omega$, and the ground-truth reward is: $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)\;,$ where $\psi(y),\psi(y,\omega)\in\mathbb{R}^{d}$, $\textbf{r}_{\textup{sparse}}\in\mathbb{R}^{d}$, $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, and $k\ll d$.
@@
where $\textbf{r}_{\textup{sparse}}$ gets entangled with $\pi_{{\mathsf{ref}}}$ in $\theta_{p,0}^{\star}$. Note that if $\log\mathbb{E}_{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}\exp(\psi(y,\omega)_{1})$ can be mapped to certain complicated non-linear function of $\psi(y)$, then the policy model is mis-specified while the reward model is not, as in Condition 2. And even without model mis-specification, we can show that an implicit representation gap can still arise due to statistical inefficiencies.
Theorem 10 (Efficient Sparse Recovery).
For DTSP task, and under mild assumptions, there exists an environment, s.t. by estimating from $n$ samples with $\theta_{1}=e_{1}$ constraint, the data-dependent estimation error of the reward model can be reduced to $\tilde{\mathcal{O}}(k\log d/n)$ using an $\ell_{0}$-constrained estimator, and can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator.
Theorem 11 (Separation).
For DTSP task, and under mild assumptions, there exists an environment, s.t. by estimating from $n$ samples with $\theta_{1}=e_{1}$ constraint, the data-dependent estimation error of the reward model can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using an $\ell_{1}$-regularized estimator, while the estimation error of the policy model is lower bounded by $\Omega(d/n)$.
@@
Benefiting from fixing optimal $\theta_{1}$, which is relatively easier to estimate, we can reduce the dual-token prediction problem to one-token prediction, where $\theta^{\star}_{r,0}$ is sparse while $\theta^{\star}_{p,0}$ is dense. Pure reward learning can thus reduce the estimation error to $\mathcal{O}(1)$ using $\mathcal{O}\left(k\log d\right)$ samples instead of $\mathcal{O}(d)$ samples. Detailed proofs deferred to Section C.10.
Takeaways. This section shows that the policy model is more prone to model mis-specification than the reward model, under a realistic setting, and thus Conditions 1 and 2 are more prevalent than Condition 3. And even without model mis-specification, the estimation error can also induce an implicit representation gap. From the perspective of sparse recovery, we can see that the surrogate reward model could suffer from severe statistical inefficiency, while the reward model can recover the reward structure very efficiently, even with the same model scale. Given the insight that real-world rewards are often sparse and simple (Yao et al., 2025), we can infer that the reward model’s quality typically surpasses that of the surrogate reward model. This further explains why two-stage RLHF is empirically observed to outperform DPO (Ivison et al., 2024; Xu et al., 2024b).

### v2

@@ 3.2 Policy Model Mis-specification
In this case, RLHF has a structural advantage: it can recover the exact reward and then compute the best possible policy within $\Pi$. In contrast, DPO bypasses reward modeling and directly learns a policy from preferences, which may lead to sub-optimal behavior due to mismatches between preference-based objectives and reward-based value functions. The following proposition provides a concrete example where DPO fails to recover the best achievable policy, even under exact optimization. Proof deferred to Section B.3.
@@
Furthermore, we show that online DPO cannot close this gap, even when equipped with PILAF sampler. A numerical proof is deferred to Section B.8.
@@
Remark 3.
Our key insight is that a strict performance gap between RLHF and DPO can exist under policy model mis-specification, and importantly, even sophisticated samplers like PILAF may fail to close the gap, an important nuance that, to our knowledge, has been overlooked in prior studies.
@@ 3.3 Reward Model Mis-specification
We now consider the setting where the ground-truth reward function $r^{\star}$ is not realizable by the reward model class $\mathcal{F}$, while the optimal policy $\pi^{\star}$ lies within the policy class $\Pi$. As discussed in Swamy et al. (2024), two-stage RLHF can only lose information during reward learning, which will be highlighted under reward model mis-specification.
@@
In this setting, RLHF is vulnerable to reward mis-specification: the learned mis-specified reward model $r_{\textup{RLHF}}$ could significantly deviate from the ground-truth reward $r^{\star}$, causing the subsequent policy optimization to yield a sub-optimal solution even though $\pi^{\star}\in\Pi$. Conversely, DPO has a clear advantage: it can directly fit a policy to the observed preference data and thus recover $\pi^{\star}$ without incurring reward modeling error. Proof deferred to Section B.4.
@@
Observation under token-level parameterization. To assess the practicality of Condition 3 for auto-regressive language models, we specialize our general bandit model to the token-level parameterization. In this setting, the optimal policy admits the closed-form characterization of Rafailov et al. (2024), which we restate with an explicit separation between $\pi_{{\mathsf{ref}}}$ and the $q^{\star}$ function (see Section B.11 for details):
$\pi^{\star}(y_{t}|x,y_{0\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|x,y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|x,y_{0\ldots t-1})}{\beta}\right)\;,$ (1)
where the $q^{\star}$ function is determined in a recursive way:
[…]
This observation shows that while the reward model in RLHF only needs to approximate $r^{\star}$, the policy model in DPO must capture the token-level $q^{\star}$ function, which recursively entangles the reward signal with the base model $\pi_{{\mathsf{ref}}}$. As a result, the policy model faces a substantially more demanding learning objective, making it more prone to mis-specification and suggesting that the “weak reward, strong policy model” regime may be less common in practice.
@@ 3.4 Double Model Mis-specification
To enable a fine-grained comparison between RLHF and DPO under this double mis-specified regime, we introduce the surrogate reward model class induced by the policy class as $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\theta\in\mathbb{R}^{d_{P}},\hat{r}_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{{\mathsf{ref}}}(y|x)},\forall x\in\mathcal{X},y\in\mathcal{Y}\}$. Pairwise preferences depend only on reward differences, so reward functions are equivalent if they differ by a constant. We compare the expressiveness of the original reward model class $\mathcal{F}$ and the surrogate class $\mathcal{F}_{\Pi}$, modulo constant shifts, and analyze three representative regimes characterizing their relative capacities:
@@
Analysis of the isomorphic case. Condition 5 indicates the scenario when the reward model class and policy model class are isomorphic—meaning there exists a shared parameterization or a deterministic mapping between rewards and policies. This structure allows us to directly compare RLHF and DPO when both operate under the same representational constraints, and to investigate whether bypassing reward modeling, as in DPO, provides any advantage. In RLHF, reward learning is decoupled from the current policy, and thus lacks access to its distributional information; while DPO can mitigate this limitation through online sampling. Therefore, RLHF under Condition 5 is comparable to offline DPO, but could underperform online DPO. Proofs deferred to Sections B.5 and B.9.
@@
On the other hand, under Conditions 6 and 7, either method may outperform the other depending on the environment. Proofs deferred to Sections B.6 and B.7.
@@
This result implies a general principle: the performance gap is reflected in the quality gap between the (surrogate) reward models: $r_{\textup{RLHF}}$ and $\hat{r}_{\textup{DPO}}$. Better reward learning yields higher expected value.
As revealed in Appendix C of Ouyang et al. (2022) and Section 3.3 of Swamy et al. (2025), it is uncommon to deploy a reward model with a larger scale than the policy model. And thus to ensure practical relevance, we focus on the regime $\mathcal{F}\subseteq\mathcal{F}_{\Pi}$, and pose the following relevant question:
@@
As an answer to this question, we note that in the context of preference learning, the reward model quality can be measured using an $\ell_{2}$ distance of pairwise difference, derived by simple calculations:
$V_{r_{\phi}}^{\pi_{\theta^{\star}(r_{\phi})}}\textit{ can be measured by }-\mathbb{E}_{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta^{\star}(r_{\phi})}\right)}\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(r_{\phi}(y)-r_{\phi}(y^{\prime}))\right]^{2}\;,$ (3)
where $\pi_{\theta^{\star}(r_{\phi})}:=\operatorname{argmax}_{\pi\in\Pi}\;V_{r_{\phi}}^{\pi}$ and we omit prompts for simplicity. Detailed calculations and further discussions deferred to Appendix A. Using this metric, we can further establish a separation.
Concluding remarks. Although we adopt relatively simple techniques, these results can provide valuable insights for the fundamental differences between RLHF and DPO. In the next section, we demonstrate that these insights extend naturally to more practical and realistic scenarios.
4 Approximate Optimization: Performance Gap Induced by Statistical Efficiency Differences in Reward Learning
With limited preference data, we are not able to directly compute exact solutions, and thus obtain weaker reward models and policy models due to estimation error. This scenario can be viewed as inducing an implicit model mis-specification, whose effects have been widely discussed in Condition 4 of Section 3. And since we can only lose information in reward learning (Swamy et al., 2025), Equation 2 still holds asymptotically with on-policy sampling. Thus by assuming $F\subseteq F_{\Pi}$, we only need to compare the reward model quality measure shown in Equation 3. We adopt an empirical proxy for this notion, data-induced semi-norm (details in Definition 2 in Section B.10, see also Zhu et al. (2023)): $\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\phi}(y_{w}^{(i)})-r_{\phi}(y_{l}^{(i)}))\right]^{2}$, where $\mathcal{D}=\{(y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n}$ is an empirical preference dataset and we omit prompts from now on.
@@
Difference in token-level linear parameterization. In this section, to rigorously establish a separation, we focus on a specific token-level linear parameterization, which is a special case of the general bandit model; therefore, previous results continue to hold. The common reward model shares the same architecture with LM but replaces the last layer with a linear head, i.e., it takes the whole prompt-response pair as the input and predicts one value. Therefore, if we view the last-layer hidden state as the feature vector, it is natural to assume the reward model to be parameterized as a linear MDP model [^1] It is also common to assume the reward model to be a linear bandit model (Zhu et al., 2023), while the stronger linear MDP model assumption here is for fair comparison with the following policy model.: $r_{\theta_{r}}(y)=\beta\sum_{t=0}^{n}{\theta_{r,t}}^{\top}\psi(y_{0\ldots t})$, where $\theta,\psi(y)\in\mathbb{R}^{d}$, and n=|y|1n=|y|-1. While for the policy model, one needs to go through the softmax results of all tokens and multiply them [^2] Our parameterization assumption on the token-level policy model is different from Razin et al. (2025a), which utilizes a form of token matrix, since we intend to ensure that $d_{P}=d_{R}$.:
$\pi_{\theta_{p}}(y)=\prod_{t=0}^{n}\pi_{\theta_{p,t}}(y_{t}|y_{0\ldots t-1})=\prod_{t=0}^{n}\frac{\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t}))}{\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(\theta_{p,t}^{\top}\psi(y_{0\ldots t-1},s))}\;.$
@@
Let the ground truth reward be $r^{\star}(y)=\beta\sum_{t=0}^{n}(\theta_{t}^{\star})^{\top}\psi(y_{0\ldots t})$, then the optimal solution for the reward model is $\theta_{r,t}^{\star}=\theta^{\star}_{t}$. And recall Equation 1, the optimal solution for the policy model is:
@@
$\pi_{\theta^{\star}_{p,t}}(y_{t}|y_{0\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})}{\beta}\right)\;.$
@@
Benefiting from the token-level $q^{\star}$ function, models trained in this way can simulate a process reward model to provide fine-grained information (Yuan et al., 2024; Cui et al., 2025; Shi et al., 2024; Xu et al., 2025). However, simultaneously, learning the $q^{\star}$ function sacrifices statistical efficiency due to the need to model the complicated structure. Next, we will present a concrete example to illustrate the statistical gap between pure reward learning and surrogate reward learning.
@@
Dual-token sparse prediction (DTSP) task.   The policy model is required to sequentially output two tokens, $y$ and $\omega$, and the ground-truth reward is: $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)\;,$ where $\psi(y),\psi(y,\omega)\in\mathbb{R}^{d}$, $\textbf{r}_{\textup{sparse}}\in\mathbb{R}^{d}$, $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, and $k\ll d$.
@@
where $\textbf{r}_{\textup{sparse}}$ gets entangled with $\pi_{{\mathsf{ref}}}$ in $\theta_{p,0}^{\star}$. Note that if $\log\mathbb{E}_{w\sim\pi_{{\mathsf{ref}}}(\cdot|y)}\exp(\psi(y,\omega)_{1})$ can be mapped to certain complicated non-linear function of $\psi(y)$, then the policy model is mis-specified while the reward model is not, as in Condition 2. And even without explicit model mis-specification, we can establish a separation in reward model qualities due to statistical efficiency differences.
Theorem 10 (Informal).
Under token-level linear parameterization and mild assumptions, there exists an environment for DTSP task, s.t. by estimating from a preference dataset $\mathcal{D}$ with $n$ samples under $\theta_{1}=e_{1}$ constraint, the estimation error of the reward model $\hat{\theta}_{r}$ can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator, i.e., w.p. $1-\delta$,
$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{r}}(y_{w}^{(i)})-r_{\hat{\theta}_{r}}(y_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
while the estimation error of the DPO model $\hat{\theta}_{p}$ is lower bounded by $\Omega(d/n)$:
$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
@@
By fixing the optimal $\theta_{1}$, which is relatively easier to estimate, we can reduce the dual-token prediction problem to a single-token prediction problem, where $\theta^{\star}_{r,0}$ is sparse while $\theta^{\star}_{p,0}$ is dense. Leveraging the results of Yao et al. (2025) then yields the separation. Formal statement and detailed proof deferred to Section B.10.
Concluding remarks. This section shows that the estimation error can also induce an implicit model mis-specification. From the perspective of sparse recovery, we can see that the DPO could suffer from severe statistical inefficiency compared with pure reward learning, even with the same model scale. Although our task construction is specific, it reveals a general phenomenon: DPO can distort the intrinsic structure of the true reward function. Given the insight that real-world rewards are often sparse and simple (Yao et al., 2025), we can infer that the reward model’s quality typically surpasses that of the surrogate reward model. This further explains why two-stage RLHF is empirically observed to outperform DPO (Ivison et al., 2024; Xu et al., 2024b).